# Understanding the Uncertainty of LLM Explanations: A Perspective Based on Reasoning Topology

### v1

@@ 4.2. Topology-enabled Reasoning Quantification
Our framework enables multidimensional uncertainty analysis through structural and semantic examination of reasoning topologies. Given a query $x^{q}$, the model $M$ will be asked $L$ times for explanation elicitation, on which $L$ reasoning topologies $\{\mathcal{G}^{q}_{i}\}_{i=1}^{L}$ will be generated from the previous step. We can measure the consistencies of $\{\mathcal{G}^{q}_{i}\}_{i=1}^{L}$, where both graph structure for reasoning topology, and the embeddings of node and edge sentences for semantics will be considered.
4.2.1. LLM Reasoning Uncertainty Based on Graph Edit Distance
We quantify structural uncertainty through comparative analysis of multiple reasoning topologies $\{\mathcal{G}_{i}\}_{i=1}^{L}$ generated for the same query $x^{q}$. Traditional graph comparison method (Bai et al., 2020) focuses on matching ‘sets’ of node embeddings in a broad sense (e.g., across various structured data domains like social networks (Huang et al., 2024) and chemistry (Guzman-Pando et al., 2024)), but we would expect to quantify based on the reasoning steps, which requires a more fine-grained design of distance measure. To tackle the above issue, we first use context-aware embeddings for semantic encoding, and then design a fine-grained, reasoning-step based Graph Edit Distance (GED). Specifically, we compare the reasoning structure by jointly considering semantic similarity and structural alignment in a three-step process:
@@ 4.2.1. LLM Reasoning Uncertainty Based on Graph Edit Distance
Step1: Semantic Embedding. In order to measure semantic meanings of reasoning steps, for each graph $\mathcal{G}\in\{\mathcal{G}^{q}_{i}\}_{i=1}^{L}$ we employ an embedding function $\mathcal{L}$ to encode the representation of nodes and edges in graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Since each node $v\in\mathcal{V}$ and edge $e\in\mathcal{E}$ serves as a textual description, we can derive contextual embeddings:
(14) $\mathbf{h}_{v}=\mathcal{L}(v),\quad\mathbf{h}_{e}=\mathcal{L}(e),\quad\forall v\in\mathcal{V},e\in\mathcal{E}$
@@
In this paper, we use BERT as our embedding function $\mathcal{L}$ but other embeddings could also be used for different domain contexts, e.g, (Rasmy et al., 2021) for medical text. This step encodes the semantics of nodes and edges while preserving the logical structure of the reasoning process.
@@
Step2: Reasoning Topology Distance. In our setting, we have $L$ reasoning structures $\{\mathcal{G}^{q}_{i}\}_{i=1}^{L}$ generated. To measure the pairwise distance of two reasoning structures $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, inspired by the concept of graph edit distance (Gao et al., 2010), we use the minimum transformations required to align the two graphs to quantify their pairwise distance.
@@
A. Substitution Costs: For two corresponding nodes in different reasoning topology graphs $\mathcal{G}_{1}=(\mathcal{V}_{1},\mathcal{E}_{1})$ and $\mathcal{G}_{2}=(\mathcal{V}_{2},\mathcal{E}_{2})$ we define the semantic substitution cost based on $v_{i}\in\mathcal{V}_{1}$ and $e_{k}\in\mathcal{E}_{1}$:
@@
(15) $c(v_{i},v_{j},e_{k},e_{m})=\begin{cases}1-\cos(\mathbf{h}_{i}^{v},\mathbf{h}_{j}^{v}),&\text{node substitute}\\ 1-\cos(\mathbf{h}_{k}^{e},\mathbf{h}_{m}^{e}),&\text{edge substitute}\end{cases}$
@@
where cosine similarity measures the semantic alignment of reasoning steps and elicitation questions, this can capture the difference from direct meaning: either given the similar elicitation (edge sub-question), the sub-response (nodes) are different - there might exist an incorrect answer, or different edges that lead to the similar response - there might be a jumping step.
B. Deletion/Insertion Costs: The cost of deleting a node or an edge is computed based on its average similarity to other nodes or edges within the same reasoning topology $\mathcal{G}$ (Serratosa, 2021; Bai et al., 2018). For the two reasoning graphs, we compute the deletion cost for a node $v_{i}\in\mathcal{V}_{1}$ or an edge $e_{k}\in\mathcal{E}_{1}$ with respect to the graph $\mathcal{G}_{1}$, as follows:
(16) $c_{\text{del.}}(v_{i},v_{j},e_{k},e_{m})=\begin{cases}1-\frac{1}{|\mathcal{V}_{1}|-1}\sum\limits_{\begin{subarray}{c}v_{j}\in\mathcal{V}_{1},\\ v_{j}\neq v_{i}\end{subarray}}\cos(\mathbf{h}_{i}^{v},\mathbf{h}_{j}^{v}),&\text{node delete},\\[6.0pt] 1-\frac{1}{|\mathcal{E}_{1}|-1}\sum\limits_{\begin{subarray}{c}e_{m}\in\mathcal{E}_{1},\\ e_{m}\neq e_{k}\end{subarray}}\cos(\mathbf{h}_{k}^{e},\mathbf{h}_{m}^{e}),&\text{edge delete}.\end{cases}$
where ()($\cdot$)( ⋅ ) is the same as in Eq. 15, $\sum\limits_{\begin{subarray}{c}v_{j}\in\mathcal{V}_{1},v_{j}\neq v_{i}\end{subarray}}\cos(\mathbf{h}_{i}^{v},\mathbf{h}_{j}^{v})$ shows the semantic connectivity from node $v_{i}$ to other nodes in the topology, if the embedding meaning is closer to other nodes, the value is larger, and then the value is normalized by the number of remaining nodes $|\mathcal{V}_{1}|-1$, and subtracted from 1 to compute the deletion cost. Highly similar nodes or edges in $\mathcal{G}_{1}$ (e.g., redundant sub-questions) will have lower deletion costs, as their removal minimally impacts the reasoning flow. Conversely, unique or critical nodes and edges (e.g., important conclusions or key transitions) will incur higher deletion costs due to their significant role in maintaining reasoning integrity and structural coherence.
Step3: Graph Distance for Reasoning Uncertainty. Based on the above two steps, we can derive the overall graph edit cost in joint consideration of semantic meaning and topology variance as:
(17) $\text{GED}(\mathcal{G}_{1},\mathcal{G}_{2})=C_{\text{sub.}}(\mathcal{P})+C_{\text{del.}}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{P})$
@@
where $\mathcal{P}$ represents the optimal matchings for sub-questions (edges $\mathcal{P}_{e}$) and sub-responses (nodes $\mathcal{P}_{v}$), computed using the Hungarian algorithm (Mills-Tettey et al., 2007). The term $C_{\text{sub.}}(\mathcal{P})$ accounts for the total substitution costs over the two graphs, and $C_{\text{del.}}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{P})$ captures the total deletion costs for nodes and edges over the two graphs, details are explained in the Appendix. E.1. So we can calculate the minimal total cost of transformations by finding:
(18) $\text{GED}_{m}(\mathcal{G}_{1},\mathcal{G}_{2})=\min_{\mathcal{P}}\text{GED}(\mathcal{G}_{1},\mathcal{G}_{2})$
A higher GED implies a higher difference in the reasoning phase by considering both embedding and structures.
We use this computed reasoning distance to construct a distance matrix across multiple reasoning structures. Given a set of reasoning topologies $\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{n}\}$, we compute pairwise distances using the GED-based similarity measure: $d_{ij}=\text{GED}_{m}(\mathcal{G}_{i},\mathcal{G}_{j})$, which then forms the overall distance matrix between $k$ reasoning topologies
(19) $D^{\mathcal{G}_{n}}=[d_{ij}]=[\text{GED}_{m}(\mathcal{G}_{i},\mathcal{G}_{j})]_{n\times n}$
where each entry $d_{ij}$ quantifies the structural and semantic difference between the reasoning processes in $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$. Here, we now can resort to the UQ measure $\mathcal{U}(\cdot)$ to the variance of the distances in $D^{\mathcal{G}_{n}}$, which reflects the inconsistency or stability of the model’s reasoning behavior. Combining Eq. 3, we have the uncertainty score over a query $x^{q}$ as:
(20) $\mathcal{U}_{\text{struct}}(x^{q})=\text{Var}(D^{\mathcal{G}_{n}}),$
@@
where $D^{\mathcal{G}_{n}}$ is the pairwise distance matrix over the set of reasoning topologies $\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{n}\}$. The function $\mathcal{U}(\cdot)$ = $\text{Var}(\cdot)$ computes the overall variance of all pairwise distances. A higher variance indicates greater inconsistency in the model’s reasoning, suggesting that the LLM generates significantly different structures across multiple responses to the same query.
4.2.2. LLM Reasoning Redundancy Measure
@@
It is known that the LLM’s reasoning efficiency varies based on the problem type and model weights (Plaat et al., 2024), and the reasoning topology provides a good reference to understand the efficiency by analyzing the detailed steps. We find that LLMs do not necessarily rely on all of the nodes from its reasoning topology for the final conclusion drawing, which means that, some of the sub-steps do not contribute to solving a problem and it causes the efficiency decrease. Here, we propose a way of measurement named ‘Reasoning Redundancy’. Reflect the reasoning steps in Eq. 13: $S=\{[v_{i},v_{j},e_{ij}]\mid v_{i},v_{j}\in\mathcal{V},e_{ij}\in\mathcal{E}\}$, we aim to measure the redundancy based on the valid path constructed by the steps.
@@
$v_{k}\notin\bigcup_{[v_{i},v_{j},e_{ij}]\in\mathcal{P}_{\text{valid}}}\{v_{i},v_{j}\},$
where $\mathcal{P}_{\text{valid}}$ represents the set of all valid paths contributing to the final conclusion.
We have designed detailed criteria for efficient searching as shown in Appendix E.2, then we perform the searching for $\mathcal{P}_{valid}$ valid paths and ‘Redundancy Rate’ using traversal algorithm (DFS), then we have the redundancy rate of the reasoning process for $a^{e}_{i}$ as:
(21) $r_{\text{redun.}}(a^{e}_{i})=\frac{|\mathcal{V}_{\text{redundant}}|}{|\mathcal{V}|}$
where the $|\mathcal{V}|$ is the total number of nodes in the reasoning topology, and $|\mathcal{V}_{\text{redundant}}|$ is the number of redundant nodes.
4.2.3. A Recipe in Practice
The above-mentioned quantification methods have different properties in potential usage. The Topology-UQ is mainly proposed to evaluate the trustworthiness of LLM responses, make pair-wise reasoning path comparisons, or conduct completeness checks like what is strengthened in the AI4education domain (Ng et al., 2024), where the teaching is focusing on the logic steps rather than final answers. The reasoning Redundancy measure can be used to conduct a reasoning-efficiency check. It helps to identify the non-necessary discussions for a problem and guide the LLM for improvement by probing and marking a more concise but correct path because high-quality answering should not only focus on correctness but also involve solving efficiency.
Methods Application
Topology-UQ ①Trustworthy, ②Path similarity ③Completeness
Redundancy ①Efficiency check, ②Guide improvement
Table 1. The usage of two quantification measures.
@@ 4.2.3. A Recipe in Practice
PCC SRC KR PCC SRC KR PCC SRC KR PCC SRC KR PCC SRC KR
@@
Table 2. Comparison of our methods with different baselines on various datasets and large language models. The results show that our results consistently outperform the baseline methods. The value reflects the correlation between the uncertainty and the ground truth faithfulness, the more faithful an LLM performs, the less uncertainty it should be, so by the UQ methods, we would expect a negative correlation for a well-performed UQ method.
5. Experimental Study
In this section, we conducted extensive experiments covering 5 LLMs: GPT4o-mini, DeepSeek-R1 (distilled version on llama3-70b), Phi4, Llama3-8b, Llama3-70b (v3.3), 3 challenging datasets and using a total of 5 methods. We use the proposed method for topology-based LLM explanation analysis, then we try to answer the following research questions.
  • RQ11: Can the proposed method reveal the actual uncertainty in LLM’s reasoning?
  • RQ2: How do LLMs perform in the proposed redundancy measure?
5.1. Experiment Settings
In this section, we introduce the experiment settings including the dataset adopted for analysis and baselines used for comparison. Besides, we also introduce the evaluation metrics for measuring the UQ methods’ performance on Natural Language Explanation tasks.
5.1.1. Dataset
In this study, to align with the research community (Mirzadeh et al., 2024; Toshniwal et al., 2024; Zhang et al., 2024), we utilized widely adopted datasets, including GSM8k (Cobbe et al., 2021) and BoolQ (Clark et al., 2019), which require complex reasoning rather than simple question-answering. These benchmarks assess LLM’s ability to perform multi-step inference and logical reasoning. Besides, we also develop a new dataset, GeoQA, especially for condition-based reasoning tasks. We will explain the details of the GeoQA dataset in Appendix B and provide brief introduction of all datasets in Appendix F.1
5.1.2. Baselines
To the best of our knowledge, there are only few works focusing on the uncertainty quantification of NLE, thus, we not only included the existing method (1) Chain-of-Thought Agreement (CoTA) (Tanneru et al., 2024), but also analyze (2) Embedding distance-based UQ for NLE (Embed-UQ), (3) Entailment probability-based (Entail-UQ), and (4) NLI-logit based UQ, as our extra baselines to understand their interpretability of LLM explanations. We provide a brief introduction here, and for detailed explanation of baselines is in Appendix. D.
  • CoTA: This is designed to quantify uncertainty in LLM explanations by measuring the agreement between two Chain-of-Thought explanations. It assesses whether intermediate reasoning steps in one CoT explanation are aligned with corresponding steps in another by calculating their entailment scores.
  • Embed-UQ: It measures the uncertainty of NLE by embedding distances. Given a query $x^{q}$, collected $k$ responses and, their associated explanations $\{a^{e}_{1},a^{e}_{2},\dots,a^{e}_{k}\}$. Each explanation $a^{e}_{i}$ is embedded into a high-dimensional space, and the pairwise distances between embeddings are computed.
  • Entail-UQ: It extends Embed-UQ by replacing the embedding-based distance metric with an entailment-based similarity measure. Given explanations such as $a^{e}_{k}$, we compute an entailment similarity matrix, where the similarity between two explanations is calculated by entailment models 222off-the-shelf DeBERTa-large model.
5.1.3. Evaluation and Metrics
To evaluate the performance of uncertainty quantification methods in LLM explanation tasks, we follow the standard practice that compares uncertainty results with actual faithfulness (Tanneru et al., 2024). The ground truth faithfulness score reveals how much the model relies on its complete reasoning process, which is calculated through a strategy named ‘Early Answering’ as proposed by (Lanham et al., 2023), we provide details on how the faithfulness score is derived in Appendix F. Ideally, a UQ method is good if, for a higher faithful set, it generates lower uncertainty, and vise versa (Tanneru et al., 2024). Hence, we employ three robust statistical metrics to quantify the correlation between the derived uncertainty and faithfulness. First, we use the commonly adopted metric - Pearson Correlation Coefficient (PCC), which is to measure the linear correlation between two variables. And given the relative small amount of each bootstrap sample, we employ two extra metrics Spearman Rank Correlation (SRC) and Kendall Rank Correlation (KR), the calculation of metrics is also in Appendix.
For fair evaluation and to avoid bias in single answers, we conduct bootstrap for a given dataset $D_{test}$ and measure the correlation in each sub-set $D^{\prime}$ level between uncertainty with the same level of faithfulness score. The sub-set is cut as 20 questions with 10 responses for each question = 200 $\cdot a^{e}$ and bootstrap is conducted 1000 times on each dataset.
Figure 3. The examples of three types of reasoning patterns. As shown in the image, the first type directly reflects the correct answer in its first reasoning step, and now actually performing the reasoning, the second type is like humans, try to recall the knowledge points, and narrow down the range, and find the possible answer in a thought chain, while the third one is a combination of two, it has the steps to ask for characteristics, but it also try to reflect if it has seen this question before at the edge1, which is a direct attempt to use answer to match the question, the last step is to confirm if the memory of answer matches the requirement in the question.
5.2. Quantitative Evaluation (RQ1)
@@
In order to understand how our proposed method works in the reasoning uncertainty measure tasks for LLM explanations, we perform experiments on GSM8K, BoolQ, and GeoQa datasets. Due to the page limit, GSM8K, and BoolQ are shown in the Table. 2, our method reveals a stronger negative correlation between the derived UQ results and the groundtruth faithfulness across different statistic metrics, the results on GPT4o-mini, Llama3-70b, and DeepSeek-R1 are more convincing because they have a more stable performance on the Topology elicitation task as in Figure. 5, which, is a key step for the proposed UQ method, the performance on Phi4 and Llama3-8b not promising as ranked in the last two positions in Appendix Figure. 5. This research question result shows that our method is effective in revealing the LLM’s real faithfulness, yet it is more suitable for LLMs with good instruct-following abilities.
5.3. Redundancy Measure of LLMs (RQ2)

### v2

@@ 4.2 Topology-based Uncertainty Quantification
With the Topology Elicitation module built, we intend to understand the uncertainty from the elicited structural explanations. Given a query $x^{q}$, the model $M$ will be asked $L$ times, from which reasoning topologies $\{\mathcal{G}^{q}_{i}\}_{i=1}^{L}$ will be extracted. We propose a UQ method: Reasoning Graph Edit Distance, namely Reason-GED, to evaluate the consistencies of $\{\mathcal{G}^{q}_{i}\}_{i=1}^{L}$ from the semantics embeddings of node, edge sentences, and structural topologies.
If we apply the traditional graph distances (Bai et al., 2020; Huang et al., 2024; Guzman-Pando et al., 2024) that solely quantify the differences in structures, we will lose the node and edge meaning implied in the structure. To understand the structural variance without loss of semantic meaning, we first encode the node and edge content into semantic embeddings.
@@
Step1: Semantic Embedding. To measure semantic meanings of reasoning steps, for each graph $\mathcal{G}=(\mathcal{V},\mathcal{E})\in\{\mathcal{G}^{q}_{i}\}_{i=1}^{L}$ we employ an embedding function $\mathcal{L}$ to encode the content of nodes and edges in graph. Since the content of each node $v\in\mathcal{V}$ and edge $e\in\mathcal{E}$ is textual description, we can derive embeddings:
$\mathbf{h}^{v}=\mathcal{L}(v),\quad\mathbf{h}^{e}=\mathcal{L}(e),\quad\forall v\in\mathcal{V},e\in\mathcal{E}$ (3)
@@
This paper adopts BERT as the embedding function $\mathcal{L}$, but other embeddings could be preferred for different domain contexts, e.g, (Rasmy et al., 2021) for medical text. The encoded graph $\mathcal{G}^{\prime}$ is shown in the gradient color of ② in Figure. 2. This obtains the semantics of nodes and edges while preserving the logical structure of the reasoning process.
@@
Step2: Reasoning Distance from Topology and Semantic. In our setting, we have $L$ reasoning structures $\{\mathcal{G}^{q}_{i}\}_{i=1}^{L}$ generated. To measure the pairwise distance of two reasoning processes $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, inspired by the concept of graph edit distance (Gao et al., 2010), we use the minimum transformations (substitution, deletion, insertion) required to align the two graphs as their pairwise distance.
@@
A. Substitution Costs: For two corresponding (edge, node) pairs in different topologies $\mathcal{G}_{1}=(\mathcal{V}_{1},\mathcal{E}_{1})$ and $\mathcal{G}_{2}=(\mathcal{V}_{2},\mathcal{E}_{2})$ we define the semantic substitution cost based on $(v_{i},e_{i})\in\mathcal{V}_{1},\mathcal{E}_{1}$ and $(v_{j},e_{j})\in\mathcal{V}_{2},\mathcal{E}_{2}$:
@@
$c\{(e_{i},v_{i}),(e_{j},v_{j})\}=\begin{cases}1-\cos(\mathcal{L}(e_{i}),\mathcal{L}(e_{j}))=1-\cos(\mathbf{h}_{j}^{e},\mathbf{h}_{j}^{e}),&\text{edge substitute}\\ 1-\cos(\mathcal{L}(v_{i}),\mathcal{L}(v_{j}))=1-\cos(\mathbf{h}_{i}^{v},\mathbf{h}_{j}^{v}),&\text{node substitute}\end{cases}$ (4)
@@
where cosine similarity measures the semantic alignment of two reasoning edges and nodes. If $c\{\cdot,\cdot\}$ is high, this indicates: either given a similar edge (sub-question), the nodes (sub-response) are different - there might exist an incorrect sub-answer, or difference starts with edges, which follow up with different responses - there is a different reasoning path or jumping step. The examples for these cases are shown in Appendix C.1.
B. Deletion/Insertion Costs: In order to decide which node/edge to delete or insert to align two reasoning topologies. We define the deletion and insertion costs. For a node $v_{i}\in\mathcal{V}_{1}$ in graph $\mathcal{G}_{1}=(\mathcal{V}_{1},\mathcal{E}_{1})$ relative to target graph $\mathcal{G}_{2}=(\mathcal{V}_{2},\mathcal{E}_{2})$:
$c_{\text{del.}}(v_{i})=\frac{1}{2}\{\underbrace{\max_{v_{j}\in\mathcal{V}_{2}}\cos\big{(}\mathcal{L}(v_{i}),\mathcal{L}(v_{j})\big{)}}_{\text{Cross-graph Matching Term}}\;+\;\underbrace{1-\frac{1}{|\mathcal{V}_{1}|-1}\sum_{\begin{subarray}{c}v_{k}\in\mathcal{V}_{1}\\ v_{k}\neq v_{i}\end{subarray}}\cos\big{(}\mathcal{L}(v_{i}),\mathcal{L}(v_{k})\big{)}}_{\text{Internal Uniqueness Term}}\}.$ (5)
The Eq. 5 jointly considers the cross-graph matching cost and the internal uniqueness cost when deleting node $v_{i}$. In the cross-graph matching part, if $v_{i}$ has a very similar counterpart $v_{j}$ in $\mathcal{G}_{2}$, then deleting $v_{i}$ would lose important alignment information; thus, a high similarity here leads to a higher deletion cost. In the uniqueness part, if $v_{i}$ is very similar (i.e., redundant) to many other nodes, the average similarity will be high, making this term low. In other words, deleting a redundant node is less costly. Otherwise, if $v_{i}$ is unique, deletion is costly. Since the insert. is a reverse action of delete., they share the same equation. The above defines the cost function for a node, the same applies to edge $e_{i}$, complete Eq. 19 is in Appendix C.2.
@@
Step3: Graph Distance for Reasoning Uncertainty. Based on the above two steps, we can derive the overall graph edit cost in joint consideration of semantic meaning and topology variance as: $\text{GED}(\mathcal{G}_{1},\mathcal{G}_{2})=C_{\text{sub.}}(\mathcal{P})+C_{\text{del.}}(\mathcal{P})$, the $\mathcal{P}$ represents the optimal matchings for sub-questions (edges $\mathcal{P}_{e}$) and sub-responses (nodes $\mathcal{P}_{v}$), using the Hungarian algorithm (Mills-Tettey et al., 2007), details are explained in the Appendix. H.1. We can calculate the minimal total cost of transformations by finding:
$\text{GED}_{m}(\mathcal{G}_{1},\mathcal{G}_{2})=\min_{\mathcal{P}}\text{GED}(\mathcal{G}_{1},\mathcal{G}_{2})$ (6)
@@
Given this, a higher GED implies a higher difference in the reasoning phase by considering both embedding and structures. We use this computed reasoning distance and construct a matrix across $L$ reasoning structures $\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{l}\}$, we compute pairwise distances using: $d_{ij}=\text{GED}_{m}(\mathcal{G}_{i},\mathcal{G}_{j})$, which then forms the overall distance matrix between $L$ reasoning topologies as shown on the right of Figure. 2: $D^{\mathcal{G}_{n}}=[d_{ij}]=[\text{GED}_{m}(\mathcal{G}_{i},\mathcal{G}_{j})]_{L\times L}$. Now, we can calculate the Topo-UQ by the variance of the distances in $D^{\mathcal{G}_{n}}$, which reflects instability of the model’s reasoning behavior. Combining Eq. 2, we have the uncertainty score over a query $x^{q}$ as: $\mathcal{U}_{\text{struct}}(x^{q})=\text{Var}(D^{\mathcal{G}_{n}})$ A higher variance indicates greater inconsistency in the model’s reasoning, suggesting that the LLM generates more different structures across multiple responses to the same query.
4.3 LLM Reasoning Redundancy Measure
@@
It is known that the LLM’s reasoning efficiency varies based on the problem type and model weights (Plaat et al., 2024), and the reasoning topology provides a good reference to understand the efficiency by analyzing the valid reasoning path. We find that LLMs do not necessarily rely on all of the nodes from their reasoning topology for the final conclusion, which means some of the sub-steps do not contribute to solving a problem, and this causes the decreased efficiency. Here, we propose a measurement named ‘Reasoning Redundancy’. Reflect the reasoning steps: $S=\{[v_{i},v_{j},e_{ij}]\mid v_{i},v_{j}\in\mathcal{V},e_{ij}\in\mathcal{E}\}$, we aim to measure the redundancy based on the valid path constructed by the steps.
@@
$v_{k}\notin\bigcup_{[v_{i},v_{j},e_{ij}]\in\mathcal{P}_{\text{valid}}}\{v_{i},v_{j}\}$ (7)
where $\mathcal{P}_{\text{valid}}$ represents the set of all valid paths contributing to the final conclusion, nodeRaw and nodeResult are the initial reasoning question and result, respectively, as in Figure. 2.
Given the definition, it is feasible to detect the redundancy following the criteria in Appendix H.2, we perform the searching for $\mathcal{P}_{valid}$ valid paths and ‘Redundancy Rate’ using traversal algorithm (DFS), then we have the redundancy rate of the reasoning process for $a^{e}_{i}$:
$r_{\text{redun.}}(a^{e}_{i})=\frac{|\mathcal{V}_{\text{redundant}}|}{|\mathcal{V}|}$ (8)
where the $|\mathcal{V}|$ is the total number of nodes in the reasoning topology, and $|\mathcal{V}_{\text{redundant}}|$ is the number of redundant nodes (isolated from a valid reasoning path).
5 Experimental Study
In this section, we conducted extensive experiments covering 5 LLMs: GPT4o-mini, DeepSeek-R1 (distilled on llama3-70b), Phi4, Llama3-8b, Llama3-70b (v3.3), 3 classic datasets, and 5 UQ methods. We intend to answer the following research questions:
  • RQ1: Can Topo-UQ reveal the actual uncertainty in LLM’s reasoning?
  • RQ2: How do LLMs perform in the proposed redundancy measure?
5.1 Experiment Settings
Dataset: In this study, to align with the research community (Mirzadeh et al., 2024; Toshniwal et al., 2024; Zhang et al., 2024), we utilized widely adopted datasets, including GSM8k (Cobbe et al., 2021) and BoolQ (Clark et al., 2019), which require complex reasoning rather than simple question-answering. These benchmarks assess LLM’s ability to perform multi-step inference and logical reasoning. Besides, we also develop a new dataset, GeoQA, especially for condition-based reasoning tasks. We will explain the details of the GeoQA dataset in Appendix E and provide a brief introduction of all datasets in Appendix I.1
Baselines: To the best of our knowledge, few works focus on the uncertainty quantification of natural language explanations (NLE), thus, we included all possible baselines: (1) Chain-of-Thought Agreement (CoTA) (Tanneru et al., 2024), (2) Embedding distance-based UQ for NLE (Embed-UQ), (3) Entailment probability-based (Entail-UQ), and (4) NLI-logit based UQ, as our extra baselines to understand their interpretability of LLM explanations. We provide a brief introduction here, and for the detailed explanation of baselines, please refer to Appendix. G.
Evaluation and Metrics: To evaluate the performance of uncertainty quantification methods in LLM explanation tasks, we follow the standard practice that compares uncertainty results with actual faithfulness (Tanneru et al., 2024). The ground truth faithfulness score reveals how much the model relies on its complete reasoning process, which is calculated through a strategy named ‘Early Answering’ as proposed by (Lanham et al., 2023), We provide details on how the faithfulness score is derived in Appendix I. Ideally, a UQ method is good in NLE if, for a higher faithful set, it generates lower uncertainty, and vice versa (Tanneru et al., 2024). Hence, we employ three robust statistical metrics to quantify the correlation between the derived uncertainty and faithfulness. Including Pearson Correlation Coefficient (PCC), which is used to measure the linear correlation between two variables. Given the relatively small amount of each bootstrap sample, we employ two extra metrics, Spearman Rank Correlation (SRC) and Kendall Rank Correlation (KR); the calculation of metrics is also in the Appendix J.
For fair evaluation and to avoid bias, we conduct a bootstrap for a given dataset $D_{test}$ and measure the correlation between uncertainty and faithfulness score at each sub-set $D^{\prime}$ level. The sub-set is cut as 20 questions with 10 responses for each question = 200 $a^{e}$, and bootstrapping is conducted 1000 times on each dataset.
5.2 Quantitative Evaluation (RQ1)
@@
PCC()($\downarrow$) SRC()($\downarrow$) KR()($\downarrow$) PCC()($\downarrow$) SRC()($\downarrow$) KR()($\downarrow$) PCC()($\downarrow$) SRC()($\downarrow$) KR()($\downarrow$) PCC()($\downarrow$) SRC()($\downarrow$) KR()($\downarrow$) PCC()($\downarrow$) SRC()($\downarrow$) KR()($\downarrow$)
@@
Table 1: Comparison of our methods with different baselines on various datasets and large language models. The results show that our results consistently outperform the baseline methods. The value reflects the correlation between the uncertainty and the ground truth faithfulness; the more faithful an LLM performs, the less uncertainty it should have, so by the UQ methods, we would expect a negative correlation for a well-performed UQ method.
@@
In order to understand if Topo-UQ can reveal the uncertainty in LLMs’ reasoning, we perform experiments on GSM8K, BoolQ, and GeoQA datasets. Due to the page limit, GSM8K and BoolQ are shown in the Table. 1, GeoQA is shown in Appendix, Figure. 4. Our method reveals a stronger negative correlation between the derived UQ results and the groundtruth faithfulness across different statistic metrics, the results on GPT4o-mini, Llama3-70b, and DeepSeek-R1 are more convincing because they have a more stable performance on the Topology elicitation task as in Figure. 6, which, is a key step for the proposed UQ method, the performance on Phi4 and Llama3-8b was not promising as ranked in the last two positions in Appendix Figure. 6. This RQ result shows our method is effective in revealing the LLM’s real faithfulness, yet it is more suitable for LLMs with good instruction-following abilities.
5.3 Redundancy Measure of LLMs (RQ2)
Figure 3: The left side shows the quantification of the LLM’s reasoning redundancy on three datasets: GeoQA (Blue), GSM8K (Yellow), BoolQ (Purple); The right side shows the clustering result of the reasoning topologies.